If the inverse demand curve of profit maximizing monopolist is given as P =1200 − 2Q , and cost function as 

 C = Q3 − 61.25Q2+1528.5Q + 2000, find equilibrium output level, monopolist price, and profit.


1
Expert's answer
2021-07-27T15:15:02-0400

"P=1200-2Q"

"TR=1200Q-2Q^2"

"TC=Q^3-61.25Q^2+1528.5Q+2000"


"MR=\frac {\delta TR}{\delta Q}=1200-4Q"


"MC=\frac {\delta TC}{\delta Q}= 3Q^2-122.5Q+1528.5"

(1) To find the equilibrium output

Equate MR=MC:

"1200-4Q=3Q^2-122.5Q+1528.5"

"3Q^2-118.5Q+328.5" =0 …(equation 1)

solving equation 1 quadratically,

"Q=\frac {118.5+/- ([118.5]^2-[4\times 3\times 328.5])^\frac{1}{2}}{2\times3}"


"Q=37"

Equilibrium output level= 37 .

The profit maximizing quantity= 37

(2) To find the monopolistic price,

Substitute the value of Q in the demand curve equation:

"P=1200-2(37)"

"P=1126"

The monopolistic price is 1126.

(3) To find profit , we work out the difference between Total Revenue(TR) and Total Cost(TC)

"TC=Q^3-61.25Q^2+1528.5Q+2000"

substituting Q=37;

"TC=25,356.25"

"TR=1200Q-2Q^2"

substituting Q =37;

"TR=41,662"

"\therefore" Profit = Total Revenue- Total Cost

= "41662-25356.25"

"=16,305.75"